OpenStudy (anonymous):

One dimension of a rectangular map is 8 inches and the perimeter is 40 inches. What is the other dimension of the map? A. 14 inches B. 12 inches C. 4 inches D. 2 inches

OpenStudy (schrodingers_cat):

2(8) + 2(l) = 40 l = (40 - 16)/2 = 12 Hope this helps :)

OpenStudy (anonymous):

it does thanks :)